Testing Hyperref

Sebastian Rahtz April 1997

2

Contents

LIST OF FIGURES  5

1 First part 7 1.1 Test section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.2 Section One — cats . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

II – 1 one.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 II – 2 one.2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

1.3 Section Two — TEX is a dog . . . . . . . . . . . . . . . . . . . . . . 8 III – 1 two.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 III – 2 two.2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2 Second part 11

2.1 Section Three — Camels . . . . . . . . . . . . . . . . . . . . . . . . 11 I – 1 three.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 I – 2 three.2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

2.2 Section Four — Butterflies . . . . . . . . . . . . . . . . . . . . . . . 12 II – 1 four.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 II – 2 four.2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

2.3 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 III – 1 subsec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 III – 2 subsec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

2.4 two . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 IV – 1 Subsection 2 . . . . . . . . . . . . . . . . . . . . . . . . . . 15 IV – 2 Subsection 3 . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2.5 three . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4 CONTENTS

List of Figures

2.1 A cat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 6 Title

Chapter 1

First part, leading to 1.1 next

1.1 Our L^{A}TEX test section (leading to 1.2) for 100% of

Æhorrid X[Y]Z things, like 42

and so see section 1.1.

1.2 Section One — cats

see section 2.1 on page 11 about cats and cite [1]

II – 1 one.1

some text with a footnote^{1} and a reference to a long table, 2.1

II – 2 one.2

dogs 8 CHAPTER 1. FIRST PART

1WISH UPON A STAR

1.3 Section Two — TEX is a dog

III – 1 two.1 III – 2 two.2

1.3. SECTION TWO — TEX IS A DOG 9 cite [1] again.

10 CHAPTER 1. FIRST PART

Chapter 2

Second part

2.1 Section Three — Camels

see section 1.2

I – 1 three.1

some text with a footnote^{1}

I – 2 three.2

1OVER THE RAINBOW

12 CHAPTER 2. SECOND PART

2.2 Section Four — Butterflies and so on

II – 1 four.1 II – 2 four.2

camels Refer to with these words 2.3. INTRODUCTION 13

2.3 Introduction

III – 1 subsec III – 2 subsec

Define a marker here while this one is a PostScript picturr acting as marker:

2.4. TWO 15

2.4 two

IV – 1 Subsection 2 IV – 2 Subsection 3

xxxx Test picture xxxxx xxxxx

Figure 2.1: A cat

2.5 three

This is a reference to section 1 (2.3), subsection 1.2 (III – 2) and section 2 (2.4). References to [1, Dallas 1992].

¡¡where is equation 2.1¿¿

and does a link to a url like http://www.tug.org work?

zzzz + b  (2.1)

d − e  (2.2)
y = z  (2.3)
g = h  (2.4)

(2.5)

We need some lists:

1. oranges

2. lemons

3. beer (a) Samuel Smiths (b) Labatts

see sec1: 2.3 sec2: 2.4 eq1: 2.1 fig1: 2.1 and cite [1] again.

18 CHAPTER 2. SECOND PART

Table 2.1: A test long table

a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c

2.5. THREE 19

a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c
a a b c

a

Does this point to the second equation? Does anything point to the eqnarray (2.3)?

20 CHAPTER 2. SECOND PART

Bibliography

[1] Barceló, J. 1992. Programming an intelligent database in archaeology. In Computer Applications and Quantitative Methods in Archaeology 1991, Lock, G. & J. Moffett (eds), 21–28, Oxford: British Archaeological Reports. 7, 9, 16, 17

[Dallas 1992] Dallas, C. J. 1992. Syntax and semantics of figurative art: a formal approach. In Archaeology and the Information Age, Reilly, P. & S. Rahtz (eds), chapter 16, London: Routledge. 16

[Stankovic 1988] J. Stankovic, “Misconceptions about real-time computing: a serious problem for next-generation systems,” Computer, vol. 21, no. 10, pp. 10–19, Oct. 1988.

22 BIBLIOGRAPHY An index entry for gnus BIBLIOGRAPHY 23

An index entry for gnus 24 BIBLIOGRAPHY

An index entry for gnus BIBLIOGRAPHY 25

An index entry for gnus 26 BIBLIOGRAPHY

An index entry for gnus BIBLIOGRAPHY 27

An index entry for gnus 28 BIBLIOGRAPHY

An appendix — the Index

29

Index

animals

cats, 7 beer, 17 cats, 19 gnus, 22–27 lemons, 17 oranges, 17 PS pictures, 14 rainbows, 11 tables

long longtables, 7